# Auditing a Dutch Public Sector Risk Profiling Algorithm Using an Unsupervised Bias Detection Tool

### v1

@@ 1 Introduction
In the period 2012-2023, a simple rule-based risk profiling algorithm of the Dutch Executive Agency for Education (DUO) contributed to indirect discrimination in a control process to check whether students were (un)duly allocated a college grant [6]. A rule-based algorithm, consisting of three apparently neutral characteristics of students (type of education, age, distance to parents), was used to assign higher risk scores to students enrolled in vocational education, younger students, and students who lived near their parental address. The seemingly impartial risk assessment led to unequal treatment of students with a non-European migration background, remaining unnoticed for over 10 years and affecting more than 350,000 students [^1] Between 2012-2023, over 350,000 students were assigned a risk score by the algorithm. Our analysis focuses on the student population in two years: 2014 and 2019, which consists of a total of 298,882 students. [5]. Given this, in 2024 the Dutch Minister for Education, Culture and Science apologized on behalf of the Dutch government for indirect discrimination experienced by students with a non-European migration background and announced a EUR 61 million compensation plan [5]. This example of a new discriminatory algorithm scandal in the Netherlands [35] highlights that not only complex statistical methods, like machine learning and deep learning [16], but also simple rule-based algorithms pose significant risks of embedding discrimination in algorithm-supported decision-making processes. In a recent interview, the chair of the Dutch Data Protection Authority warned that discriminatory rule-based algorithms are “around every corner” in executive branches of the Dutch government [60].
@@
This type of risk profiling is prohibited by the Charter of Fundamental Rights of the European Union (EU) [3], the EU General Data Protection Regulation (GDPR) [4] and the EU AI Act [29], as well as the national Public Administration Law [2]. Limited institutional enforcement capabilities, along with a lack of awareness, have led to inadequate oversight of algorithmic-supported decision-making processes [48].
@@
To combat bias in algorithmic systems, the academic community has presented a multitude of ways to define, detect, and mitigate different notions of bias at different points along the algorithmic pipeline [47, 18, 11, 69]. Most bias mitigation methods require data on the demographic groups to prevent discriminatory bias. These demographic groups are typically protected attributes under discrimination law, such as race, gender and age [^2] The exact protected grounds are enshrined in European non-discrimination law, such as the Treaty on the Functioning of the European Union (TFEU), the Charter of Fundamental Rights of the EU, and specific directives like the Racial Equality Directive (2000/43/EC) and the Employment Equality Directive (2000/78/EC). The main protected attributes include: race and ethnicity, gender and sex, religion or belief, disability, age, sexual orientation, nationality and marital or family status. Throughout, we will refer to these attributes as demographic groups. Note that some provisions do not mention any particular grounds of discrimination, such as Art. 20 of the EU Charter. Instead using openly formulated clauses such as ”everyone is equal before the law”. [10]. However, data on demographic groups are often not available to organizations or external algorithm auditors [8, 15]. Upon request, public sector organizations can occasionally obtain aggregated statistics on demographic groups through official institutions, such as the National Office of Statistics, which have access to statistical population registers. Private organizations cannot request these data and face ethical, legal, and practical challenges when collecting data on demographic groups customers belong to  [15, 30, 29]. Unsupervised learning can aid in this challenge by detecting groups in the data for which a given notion of equal treatment is violated, avoiding the need for ex-ante specification of sensitive groups [49].
@@
In this paper, we use an unsupervised bias detection tool to audit the risk profiling algorithm of DUO on bias. For this audit, the country of birth and country of origin were provided by Statistics Netherlands [52]. These data enable us to verify whether the clusters identified by the unsupervised bias detection tool, which represent potentially unfairly treated groups, correspond to disadvantaged demographic groups. This allows for a comparison between an unsupervised learning approach and an approach using demographic data. As a bias metric, we use whether or not a student is categorized as “high risk” by DUO’s risk profiling algorithm. We find that the cluster with the highest bias returned by the unsupervised bias detection tool is characterized by students who are more often enrolled in vocational education, live close to their parental address, and are relatively more often 15-18 or 25-50 years old. This aligns with demographic groups found to be negatively affected by the risk profiling algorithm, based on aggregated statistics regarding students’ non-European migration background.
@@
    Advancing methodology for unsupervised bias detection: We build on existing clustering algorithms and strengthen their methodology. Using a simulation study, we examine conditions under which the tool works best and highlight key factors to consider for reliable inferences when applying it.
@@
Overall, our paper showcases the potential of unsupervised learning to aid in auditing algorithmic-supported decision-making processes via an application to a real-world use case. We argue that unsupervised bias detection serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination and unfairness in the algorithmic-supported decision-making process under review.
@@ 2.1 Challenges in Detecting and Establishing Discrimination
Under European non-discrimination law, DUO’s control process is classified as indirect discrimination as an apparently neutral practice disproportionately disadvantaged students with a non-European migration background [^3] Supra note 2 [54]. In algorithms, discrimination often arises from the proxy and correlation challenge [31]. Proxies refer to characteristics that are correlated with a demographic group, leading to disadvantaged outcomes for this group. For example, the Court of Justice of the European Union (CJEU) found that using part-time employment as a basis for differentiation in the context of occupational pension schemes constituted indirect discrimination, as over 80% of the part-time workers in Spain in 2012 were women [55]. In some cases, courts may accept that differential treatment of a demographic group is acceptable. Under EU law, indirect discrimination can be justified if it is demonstrated that a legitimate aim is being pursued, and the means of achieving that aim are appropriate and necessary [24, 25, 26, 27]. Our paper highlights how unsupervised methods can aid in the challenge of assessing indirect discrimination in algorithmic-supported decision-making processes.
@@ 2.2 Bias Detection With and Without Access to Demographic Groups
In the algorithmic fairness literature, identifying systematic differences in the treatment of individuals or groups compared to others is commonly referred to as bias detection. Most bias detection methods assume access to demographic groups such as gender or ethnicity [12, 33, 43, 40, 20]. However, in practice, these labels are often unavailable due to privacy legislation [66]. For instance, the EU’s GDPR restricts the collection of “special categories of data” (article 9) [30], which include data on ethnicity, religion, health, and sexual orientation [^4] Age and gender are not considered special categories of data. See Article 9(1) GDPR. Although, a caveat could exist [67]. The other way around, “political opinions,” “trade union membership,” “genetic,” and “biometric” data are special categories of data but are not protected by EU non-discrimination directives.. The EU AI Act introduces a potential exception to the prohibition on processing special category data: “to the extent that it is strictly necessary for the purpose of ensuring bias detection and correction” [29]. Nonetheless, ongoing legal uncertainties regarding the interplay between the AI Act and the GDPR underscore the need for methods to assess bias in algorithmic systems without relying on demographic data.
There are several alternatives to using data on demographic groups. First, researchers have presented methods to predict demographic groups, which are often referred to as proxy models or attribute classifiers [10]. For example, Bayesian Improved Surname Geocoding (BISG) [23] uses Bayes Theorem and data from the United States Census Bureau, including their surname and address, as a proxy for ethnicity or race [10]. This set of methods does require to specify the demographic group of interest beforehand. A second set of methods circumvents the problem of limited availability of protected characteristics by using a Rawlsian definition of fairness [59] that maximizes utility for the most disadvantaged group [44, 34, 17]. However, the outcome of these methods is found to not always adhere to more traditional parity-based group fairness metrics [38]. The third set of alternatives are unsupervised learning methods for fairness [50, 41]. These methods tend to modify traditional clustering algorithms to handle fairness criteria explicitly. Our work contributes to this latter category. The main benefit of this set of methods compared to proxy inferring methods is that they do not require specification of demographic groups beforehand. Thus, unsupervised learning methods are particularly helpful to examine the effect of apparently neutral provisions that might result in indirect discrimination.
@@ 2.3 Auditing Algorithmic Systems
Implementing effective auditing regimes for algorithmic systems, however, poses practical and cultural challenges. Auditors frequently encounter obstacles such as restricted data access [32], quandaries about which fairness metrics fit the context best [19] and varying opinions on what constitutes a legitimate auditor [32]. Furthermore, auditing reports are not always required to be disclosed, reducing transparency and hindering the development of public knowledge regarding auditing standards. By discussing the audit of DUO’s risk profiling algorithm, we contribute to public knowledge on best practices for auditing algorithmic systems.
@@ 3 Unsupervised Bias Detection Tool
This section introduces the theoretical background of the unsupervised bias detection tool. We use examples from our DUO case study to clarify the methodology. Figure 1 provides an overview of all steps in applying the tool.
Figure 1: Schematic overview of the steps involved in applying the unsupervised bias detection tool. The required information is a dataset, classifier, and a bias metric. Part of the dataset is used to train the Hierarchical Bias-Aware Clustering algorithm [49]. Another part of the dataset is used to test whether differences in the bias metric across clusters are statistically significant.
@@ 3.2 Clustering Algorithm
We use the Hierarchical Bias-Aware Clustering methodology from Misztal-Radecka and Indurkhya [49], hereafter referred to as “HBAC.” This is an iterative clustering algorithm that produces a partition of the original dataset to isolate potentially unfairly treated groups into distinct clusters. We adopt this algorithm because it is a standard approach for unsupervised bias detection in the realm of recommendation systems [21].
The algorithm first considers all observations in the dataset to belong to a single cluster and then iteratively splits the observations into distinct clusters. At each iteration, the algorithm identifies the cluster with the highest standard deviation of a specified bias metric, which serves as an indicator of imbalance between clusters. This cluster is then split into two subclusters. The process repeats for a predefined number of iterations, $max\_iterations$. Pseudocode for the algorithm is provided in Algorithm 1. Compared to the original algorithm from Misztal-Radecka and Indurkhya [49], we allow for clusters to be defined via $k$-modes in addition to $k$-means to allow clustering based on categorical and binary variables.
@@ 7             Add $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ to partition
The original algorithm from Misztal-Radecka and Indurkhya [49] does not specify how to assign clusters to data points that are not part of the original training dataset. We propose using the centroids of each cluster. The datapoint is assigned to the cluster whose distance from the centroid is the smallest. In the case of $k$-means and $k$-modes, the centroids are defined as respectively the mean and mode for each of the characteristics $\boldsymbol{x}_{i,k}$ [39, 37]. The Euclidean distance ($k$-means) or Hamming distance ($k$-modes) is used to determine which centroid is closer.
@@ 3.3 Minimum Number of Samples per Clusters
Algorithm 1 has been claimed to not require a specification of the number of clusters [49]. Although this is true, other hyperparameters of the algorithm such as $n_{\mathrm{min}}$ – the minimum number of samples per cluster – affect the number of clusters it finds. To determine the appropriate minimum number of samples per cluster, we use the Calinksi-Harabasz index [13] of the bias metric. This index measures the ratio of the between cluster variance over the within cluster variance. Intuitively, we want the clusters to have little variation in the bias metric within each cluster, but large variation between the clusters. This means that the data points are well separated into distinct clusters.
@@ 3.4 Testing Differences in Bias Between Clusters
After conducting the clustering, our goal is to test for each of the clusters the following null versus alternative hypothesis:
@@
For valid statistical inference, the null and alternative hypothesis must be defined before observing the data. If we conduct our statistical test on data which was also used to identify clusters, this condition is violated. Since the clusters are selected to maximize the difference in the bias metric, we will therefore tend to incorrectly conclude that there is bias when there truly is none. The broader issue of testing data-driven hypotheses—often called post-selection inference—has been studied extensively in the context of regularized regression (for an overview, see Kuchibhotla et al. [42]). A long-standing practice in that literature is to use different partitions of the data for model estimation and inference, referred to as sample splitting [46, 64]. We apply this technique by fitting the HBAC algorithm on a subset of the data and subsequently inferring the clusters and their bias on a new subset of data. Importantly, this set cannot be used to determine our hyperparameters.
When testing for multiple $K>2$ clusters, an issue is that without adjustments, the likelihood of (incorrectly) detecting a statistically significant difference under the null hypothesis increases. We address this by using a Bonferroni correction [36]. In Section 5, we illustrate that the combination of the Bonferroni correction and sample splitting prevents us from (wrongly) concluding that there is a difference in the bias metric while there is none.

### v2

@@ 1 Introduction
In the period 2012-2023, a simple rule-based risk profiling algorithm from the Dutch Executive Agency for Education (DUO) contributed to indirect discrimination in a control process to check whether students were (un)duly allocated a college grant [6]. A rule-based algorithm, consisting of three apparently neutral characteristics of students (type of education, age, and distance to parents), was used to assign higher risk scores to students enrolled in vocational education, younger students, and students who lived near their parental address. The seemingly impartial risk assessment led to unequal treatment of students with a non-European migration background, remaining unnoticed for over 10 years and affecting more than 350,000 students [^1] Between 2012-2023, over 350,000 students were assigned a risk score by the algorithm. Our analysis focuses on the student population in two years: 2014 and 2019, which consists of a total of 298,882 students. [5].
Given this, in 2024, the Dutch Minister for Education, Culture and Science apologized on behalf of the Dutch government for indirect discrimination experienced by students with a non-European migration background and announced a EUR 61 million compensation plan [5]. This example of a new discriminatory algorithm scandal in the Netherlands [35] highlights that not only complex statistical methods, like machine learning and deep learning [16], but also simple rule-based algorithms pose significant risks of embedding discrimination in algorithm-supported decision-making processes. In a recent interview, the chair of the Dutch Data Protection Authority warned that discriminatory rule-based algorithms are “around every corner” in executive branches of the Dutch government [60].
@@
This type of risk profiling is prohibited by the European Convention of Human Rights (ECHR), the Charter of Fundamental Rights of the European Union (EU) [3], the EU General Data Protection Regulation (GDPR) [4] and the EU AI Act [29], as well as the national Public Administration Law [2]. Limited institutional enforcement capabilities, along with a lack of awareness, have led to inadequate oversight of algorithmic-supported decision-making processes [48].
@@
To combat bias in algorithmic systems, the academic community has presented a multitude of ways to define, detect, and mitigate different notions of bias at different points along the algorithmic pipeline [47, 18, 11, 69]. Most bias mitigation methods require data on the demographic groups to prevent discriminatory bias. These demographic groups are typically protected attributes under discrimination law, such as race, gender and age [^2] The exact protected grounds are enshrined in European non-discrimination law, such as the Treaty on the Functioning of the European Union (TFEU), the Charter of Fundamental Rights of the EU, and specific directives like the Racial Equality Directive (2000/43/EC) and the Employment Equality Directive (2000/78/EC). The main protected attributes include: race and ethnicity, gender and sex, religion or belief, disability, age, sexual orientation, nationality, and marital or family status. Throughout, we will refer to these attributes as demographic groups. Note that some provisions do not mention any particular grounds of discrimination, such as Art. 20 of the EU Charter. Instead, using openly formulated clauses such as ”everyone is equal before the law”. [10]. However, data on demographic groups are often not available to organizations or external algorithm auditors [8, 15]. Upon request, public sector organizations can occasionally obtain aggregated statistics on demographic groups through official institutions, such as the national office of statistics, which have access to statistical population registers. Private organizations cannot request these data and face ethical, legal, and practical challenges when collecting data on demographic groups customers belong to  [15, 30, 29]. Unsupervised learning can aid in this challenge by detecting groups in the data for which a given notion of equal treatment is violated, avoiding the need for ex-ante specification of demographic or legally protected groups [49].
@@
In this paper, we use an unsupervised bias detection tool to audit the risk profiling algorithm from DUO for bias. For this audit, DUO provided data processed by their algorithm that is not publicly accessible, and Statistics Netherlands provided the aggregate data on country of birth and country of origin [52]. These data enable us to verify whether the clusters identified by the unsupervised bias detection tool, which represent potentially unfairly treated groups, correspond to disadvantaged demographic groups. This allows for a comparison between an unsupervised learning approach and an approach using demographic data. We emphasize that the clustering procedure itself does not rely on the available protected demographic data: we only use the aggregated demographic data to illustrate, post hoc, that the flagged clusters align with negatively affected groups. This analysis is therefore a validation of the clustering methodology, not a replacement for it.
In our context, the bias metric [^3] Throughout our paper, we say bias metric but this can be understood as a fairness metric as commonly used in the algorithmic fairness literature. The fairness metric reflected by our bias metric definition is demographic parity [22]. that we use requires that the probability of a student with a Dutch or European migration background being classified as ‘high risk’ by the risk profiling algorithm is equal to the probability of a student with a non-European migration background being classified as ‘high risk’. We justify the use of this metric since it assesses the adverse effect of the risk profiling algorithm on different demographic groups: a crucial aspect for investigating indirect discrimination under European non-discrimination law.
We find that the cluster with the highest bias returned by the unsupervised bias detection tool is characterized by students who are more often enrolled in vocational education, live close to their parental address, and are relatively more often 15-18 or 25-50 years old. This aligns with demographic groups found to be negatively affected by the risk profiling algorithm, based on aggregated statistics regarding students’ non-European migration background.
@@
    Advancing methodology for unsupervised bias detection: We build on existing clustering algorithms and strengthen their methodology. Through using a simulation study, we examine conditions under which the tool works best and highlight key factors to consider for reliable inferences when applying it. Our methodological improvements substantially reduce the false positive rate of the algorithm.
@@
Overall, our paper showcases the potential of unsupervised learning to aid in auditing algorithmic-supported decision-making processes via an application to a real-world use case. We argue that unsupervised bias detection serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination and unfairness in the algorithmic-supported decision-making process under review when demographic groups are unavailable.
@@ 2.1 Challenges in Detecting and Establishing Discrimination
Under European non-discrimination law, DUO’s control process is classified as indirect discrimination as an apparently neutral practice disproportionately disadvantaged students with a non-European migration background [^4] Supra note 2 [54]. In algorithms, discrimination often arises from the proxy and correlation challenge [31]. Proxies refer to characteristics that are correlated with a demographic group, leading to disadvantaged outcomes for this group. For example, the Court of Justice of the European Union (CJEU) found that using part-time employment as a basis for differentiation in the context of occupational pension schemes constituted indirect discrimination, as over 80% of the part-time workers in Spain in 2012 were women [55]. In some cases, courts may accept that differential treatment of a demographic group is acceptable. Under EU law, indirect discrimination can be justified if it is demonstrated that a legitimate aim is being pursued, and the means of achieving that aim are appropriate and necessary [24, 25, 26, 27]. Our paper highlights how unsupervised methods can aid in the challenge of assessing indirect discrimination in algorithmic-supported decision-making processes when demographic group data are not available.
@@ 2.2 Bias Detection With and Without Access to Demographic Groups
In the algorithmic fairness literature, identifying systematic differences in the treatment of individuals or groups compared to others is commonly referred to as bias detection. Most bias detection methods assume access to demographic groups such as gender or ethnicity [12, 33, 43, 40, 20, 23]. However, in practice, these labels are often unavailable due to privacy legislation [66]. For instance, the EU’s GDPR restricts the collection of “special categories of data” (article 9) [30], which include data on ethnicity, religion, health, and sexual orientation [^5] Age and gender are not considered special categories of data. See Article 9(1) GDPR. Although, a caveat could exist [67]. The other way around, “political opinions,” “trade union membership,” “genetic,” and “biometric” data are special categories of data but are not protected by EU non-discrimination directives.. The EU AI Act introduces a potential exception to the prohibition on processing special category data: “to the extent that it is strictly necessary for the purpose of ensuring bias detection and correction” [29]. Nonetheless, ongoing legal uncertainties regarding the interplay between the AI Act and the GDPR underscore the need for methods to assess bias in algorithmic systems without relying on demographic data. Algorithms can be biased even in the absence of demographic data, particularly through proxy discrimination, where a set of attributes act as a proxy for the protected attribute [10]. For example, Bayesian Improved Surname Geocoding (BISG) [23] uses Bayes Theorem and data from the United States Census Bureau, including their surname and address, as a proxy for ethnicity or race [10]. However, when information on protected characteristics is not available, the extent to which a characteristic serves as a proxy cannot be investigated and proxy detection methods are therefore not usable in these settings.
There are few alternatives to using data on demographic groups. One set of methods circumvents the problem of limited availability of protected characteristics by using a Rawlsian definition of fairness [59] that maximizes utility for the most disadvantaged group [44, 34, 17]. However, the outcome of these methods is found to not always adhere to more traditional parity-based group fairness metrics [38]. A sec set of alternatives are unsupervised learning methods for fairness [50, 41]. These methods tend to modify traditional clustering algorithms to handle fairness criteria explicitly. Our work contributes to this latter category. The main benefit of this set of methods compared to proxy inferring methods is that they do not require specification of demographic groups beforehand. Thus, unsupervised learning methods are particularly helpful to examine the effects of apparently neutral provisions that might result in indirect discrimination.
@@ 2.3 Auditing Algorithmic Systems
Implementing effective auditing regimes for algorithmic systems, however, poses practical and cultural challenges. Auditors frequently encounter obstacles such as restricted data access [32], quandaries about which fairness metrics fit the context best [19], and varying opinions on what constitutes a legitimate auditor [32]. Furthermore, auditing reports are not always required to be disclosed, reducing transparency and hindering the development of public knowledge regarding auditing standards. By discussing the audit of DUO’s risk profiling algorithm, we contribute to public knowledge on best practices for auditing algorithmic systems.
@@ 3 Unsupervised Bias Detection Tool
This section introduces the theoretical background of the unsupervised bias detection tool. We use examples from our DUO case study to clarify the methodology. We highlight that our research provides a realistic case-study of applying the tool to investigate its usefulness when one lacks protected attribute data. Our study is of both academic and practical interest since the initial algorithm is easy to apply but lacked justification, and because our data is both comprehensive and from a high-stakes policy use-case. Figure 1 provides an overview of all steps in applying the tool.
Figure 1: Schematic overview of the steps involved in applying the unsupervised bias detection tool. The required information is a dataset, a classifier, and a bias metric. Part of the dataset is used to train the Hierarchical Bias-Aware Clustering algorithm [49]. Another part of the dataset is used to test whether differences in the bias metric across clusters are statistically significant.
@@ 3.2 Clustering Algorithm
In this paper, we use the Hierarchical Bias-Aware Clustering methodology from Misztal-Radecka and Indurkhya [49], hereafter referred to as “HBAC.” HBAC is an iterative clustering algorithm that produces a partition of the original dataset to isolate potentially unfairly treated groups into distinct clusters. We use this algorithm as a starting point for the DUO audit, as it represents a standard method for unsupervised bias detection in recommendation systems [21]. Additionally, we will propose methodological enhancements to refine the algorithm for real-world AI auditing applications.
The HBAC algorithm first considers all observations in the dataset to belong to a single cluster and then iteratively splits the observations into distinct clusters. At each iteration, the algorithm identifies the cluster with the highest standard deviation of a specified bias metric, which serves as an indicator of imbalance between clusters. This cluster is then split into two subclusters. The process repeats for a predefined number of iterations, $max\_iterations$. Pseudocode for the algorithm is provided in Algorithm 1. Compared to the original algorithm from Misztal-Radecka and Indurkhya [49], we allow for clusters to be defined via $k$-modes in addition to $k$-means to allow clustering based on categorical and binary variables.
@@ 7             Add $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ to partition
The original algorithm from Misztal-Radecka and Indurkhya [49] does not specify how to assign clusters to data points that are not part of the original training dataset. We propose using the centroids of each cluster. The datapoint is assigned to the cluster whose distance from the centroid is the smallest. In the case of $k$-means and $k$-modes, the centroids are defined as, respectively, the mean and mode for each of the characteristics $\boldsymbol{x}_{i,k}$ [39, 37]. The Euclidean distance ($k$-means) or Hamming distance ($k$-modes) is used to determine which centroid is closer.
@@ 3.3 Minimum Number of Samples per Clusters
Algorithm 1 has been claimed to not require a specification of the number of clusters [49]. Although this is true, other hyperparameters of the algorithm such as $n_{\mathrm{min}}$ (the minimum number of samples per cluster) affect the number of clusters it finds. To determine the appropriate minimum number of samples per cluster, we use the Calinksi-Harabasz index [13] of the bias metric. This index measures the ratio of the between cluster variance over the within cluster variance. Intuitively, we want the clusters to have little variation in the bias metric within each cluster, but large variation between the clusters. This means that the data points are well separated into distinct clusters.
@@ 3.4 Testing Differences in Bias Between Clusters
In the original paper, no detailed methodology is outlined for testing whether there is a statistically significant difference in the bias metric between the clusters identified by the HBAC algorithm. The problem is that it is not immediately evident that the most deviating cluster, with respect to the bias metric is statistically significantly different from data points outside this cluster. To address this, we formulate a null hypothesis that assumes no difference in bias between the most deviating cluster and the rest of the dataset. Using a bootstrapped null-distribution, we test for differences between clusters, allowing for the construction of a valid statistical hypothesis test. We further suggest, for appropriate statistical testing, to apply sample splitting and adjustments for multiple hypothesis testing.
After HBAC clustering, we test the following null versus alternative hypotheses for each of the identified clusters:
@@
For valid statistical inference, the null and alternative hypotheses must be defined before observing the data. If we conduct our statistical test on data that was also used to identify clusters, this condition is violated. Since the clusters are selected to maximize the difference in the bias metric, we will therefore tend to incorrectly conclude that there is bias when there truly is none. The broader issue of testing data-driven hypotheses—often called post-selection inference—has been studied extensively in the context of regularized regression (for an overview, see Kuchibhotla et al. [42]). A long-standing practice in that literature is to use different partitions of the data for model estimation and inference, referred to as sample splitting [46, 64]. We apply this technique by fitting the HBAC algorithm on a subset of the data and subsequently inferring the clusters and their bias on a new subset of data. Importantly, this set cannot be used to determine our hyperparameters.
When testing for multiple $K>2$ clusters, an issue is that without adjustments, the likelihood of (incorrectly) detecting a statistically significant difference under the null hypothesis increases. We address this by using a Bonferroni correction [36]. In Section 5, we illustrate that the combination of the Bonferroni correction and sample splitting prevents us from (wrongly) concluding that there is a difference in the bias metric while there is none. Thus, we propose methodological improvements to the original HBAC algorithm that minimize false bias detection when using cluster analysis.